OpenStudy (yonatanbesrat):

matty jogs 9 km/hr compute matty's speed in m/s

OpenStudy (okdutchman7):

What have you figured out already?

OpenStudy (okdutchman7):

You will have to convert km/hr to m/s

OpenStudy (okdutchman7):

\[\frac{ 9km }{ hr }\times \frac{ 1000m }{ 1km }\]

OpenStudy (okdutchman7):

This converts the kilometres to metres.

OpenStudy (okdutchman7):

Do you understand?

OpenStudy (yonatanbesrat):

no what is the answer

OpenStudy (okdutchman7):

I am not allowed to just give you the answer. It violates the CoC. I can only "guide" you through the problems.

OpenStudy (okdutchman7):

Do you want my help or not?